• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen, hersman

    Log entry time 16:02:03 on August25,2003

    Entry number 113464

    keyword=Day shift summary for 25 August 2003

    Day shift summary for 25 August 2003




    finishing up kinematics 12.20 before pass change.
    Started ramp up Septum for training quench.
    9:00 Controlled access: Vince tried to change the Septum ramping speed, but the Spetum PC did not work.
    I took out the target collimators, inserted thick sieve slits, adjusted the collimators arond the sieve. Checked ice.
    Sirish went into hall to move the Compton electron detector out.
    ~12:00 back to beam permit
    14:20 Pass change completed, Beam study now, will be until 16:00.
    14:30 Controlled access: Ed and his people are working on right Dipole
    and other magnets. Patricia is checking laser optics alignment.
    15:00 Moved to pol. He3 position, rotated field to 270 degree
    15:30 transverse lasers aligned by Patricia, rotated back to 0 degree.
    16:00 Dipole is ramping back up (315 A now). Right Q1 helium level is
    near zero (it went down from 13:00-14:00). Ed and his people are trouble shooting now.